# Optimizing importance weighting in the presence of sub-population shifts

### v1

@@ 1 Introduction
Secondly, we introduce an estimation procedure for the optimal weights via a bi-level optimization problem, where the weights and the model parameters are optimized simultaneously. In the context of last-layer retraining of DNNs in the presence of sub-population shifts, we show empirically, for benchmark vision and natural language processing datasets, that existing state-of-the-art importance weighting methods (see Section 2) can be improved significantly by optimizing for the weights. We find that both the weighted average and worst-group accuracy generally improve, and that optimized weights increase robustness against the choice of hyperparameters for training. We show that the effect of optimizing weights is larger when the limited size of the training sample becomes pressing, namely in the case of a small total sample size or when the minority groups are small.
@@ 2 Related work
Sub-population shifts: it has been widely documented that ML models frequently face deteriorating generalization performance due to sub-population shifts (Cai et al., 2021; Koh et al., 2021; Yang et al., 2023). There exist different possible definitions of the groups for which the distribution shift occurs. In the case of a spurious correlation, groups are commonly defined by a combination of a label and a non-causally related attribute (Joshi et al., 2022; Makar et al., 2022). Alternatively, groups can be defined exclusively via an attribute (Martinez et al., 2021b) or, in the case of class imbalance, exclusively via a label (Liu et al., 2019). While our experiments consider spurious correlations, the methodology also applies to alternative group definitions. Generalizing to groups which do not occur in the training data (Santurkar et al., 2020) is beyond the scope of this paper.
@@
Estimation of optimal weights: a commonly noted downside of importance weighting is that it reduces the effective sample size available for model training (Martino et al., 2017; Schwartz & Stanovsky, 2022). In light of this, Cui et al. (2019) and Wen et al. (2020) suggest that the weight given to a particular label should be based on the effective sample size. Similar to our approach, Ren et al. (2018) aim to learn weights for neural networks given a validation set.
@@ 3 Problem Setting
Consider random variables $(y,\bm{x}),$ where $y\in\mathcal{Y}$ is the target variable and $\bm{x}\in\mathcal{X}$ represents the input features. Suppose we are dealing with a family of models $\left\{f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}\,|\,\bm{\theta}\in\Theta\right\}$ and we aim to find a model that minimizes $\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[\mathcal{L}(y,f_{\bm{\theta}}(\bm{x})],$ where $\mathcal{L}(\cdot,\cdot)$ is some pre-defined loss function and $\mathcal{P}_{\mathrm{te}}$ is a test distribution of interest. Moreover, we have access to a dataset $D_{\mathrm{tr}}^{n}=\{y_{{i}},\bm{x}_{i}\}^{n}_{i=1},$ which is an i.i.d. sample drawn from a training distribution $\mathcal{P}_{\mathrm{tr}}$ that is different from the test distribution. This distribution shift can be problematic, as the minimum of an empirical loss for the training data does typically not minimize the expected loss with respect to the test distribution.
@@
If we choose the weights $\bm{w}$ to be equal to the likelihood ratios $\bm{r},$ then by the law of large numbers $\mathcal{L}_{\mathrm{train}}^{n}(\bm{\theta},\bm{r})$ converges in the limit of $n$ to infinity to the expectations in Equation 2. As a consequence, $\hat{\bm{\theta}}_{n}(\bm{r})$ minimizes for sample size tending to infinity the expected loss with respect to the test distribution, $\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[\mathcal{L}(y,f_{\bm{\theta}}(\bm{x})].$ However, in practice one works with a finite amount of training data and the expectation might be poorly approximated by its sample analogue. Therefore, the optimal choice $\bm{w}^{*}_{n}$ for the weights should be defined as the one that leads to a “best” parameter estimate $\hat{\bm{\theta}}_{n}(\bm{w}),$ in the sense that it minimizes the expected loss with respect to the test distribution,
@@ 4 Theoretical analysis of linear regression
We consider a DGP $y=\bm{x}^{\prime}\bm{\beta}^{g}+\varepsilon,$ where $y\in\mathbb{R}$ is the target variable, $\bm{x}=(1,\tilde{\bm{x}}^{\prime})^{\prime}$ represents the intercept and the feature vector $\tilde{\bm{x}}\in\mathbb{R}^{d},$ and $\varepsilon\in\mathbb{R}$ is the noise. We take $\tilde{\bm{x}}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ and the noise to have zero mean, variance $\sigma^{2}$ and to be uncorrelated with $\bm{x}.$ The population consists of two groups labeled by $g\in\{0,1\}.$ Group membership determines the value of the intercept, $\bm{\beta}^{g}=(a_{g},\bm{\beta}^{\prime})^{\prime}$ for $g=0,1.$ Note that the slope coefficients are common among the two groups. The sub-population shift between training and test data is defined by a change in the marginal distribution of the group label, $p_{\mathrm{tr}}:=\mathbb{P}_{\mathrm{tr}}(g=1)\neq\mathbb{P}_{\mathrm{te}}(g=1)=:p_{\mathrm{te}}.$
@@
[…]
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}\left[(y-\bm{x}^{\prime}\hat{\bm{\beta}}_{n}(\bm{p}))^{2}\middle|D_{\mathrm{tr}}^{n}\right]\right]\approx\mathcal{B}^{2}(p_{\mathrm{te}},p,a_{1},a_{0})+\mathcal{V}(\sigma^{2},p,p_{\mathrm{tr}},n,d)+\sigma^{2},$ (8a)
@@
This is the optimal trade-off between minimizing the bias and variance. In the limit of $n\rightarrow\infty$ the variance disappears and the optimal $p$ is indeed equal to the heuristic choice $p_{\mathrm{te}}.$ However, for finite $n$ it can differ considerably. This is illustrated in Figure 1. Its impact on the expected loss is shown in Figure 5 in Appendix B.
@@ 5.1 Group-Weighted Empirical Risk Minimization
We solve the bi-level optimization problem iteratively. We find $\hat{\bm{p}}^{*}_{n}$ in Equation 10 by means of exponentiated gradient descent (Nemirovsky & Yudin, 1983; Kivinen & Warmuth, 1997), where the parameter estimates $\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})$ are kept fixed. After each weight update, we re-calibrate the parameter estimates of Equation 1. Although convergence to the optimal group weights is not guaranteed due to potential non-convexity, gradient-based optimization of hyperparameters has previously been applied successfully (Pedregosa, 2016). Details of the optimization procedure can be found in Algorithm 1.
@@ 5.3 Group Distributional Robust Optimization
where (recall that) $\mathcal{P}_{G}$ is the $G$-dimensional probability simplex. The inner expectation of Equation 13 is estimated with a validation set that is withheld from the data used to train the model parameters. The bi-level optimization procedure now consists of updating three quantities per update step: the model parameters $\bm{\theta},$ the group weights $\bm{p}$ and the loss weights $\bm{q}.$ See Algorithm 2 in Appendix D for details.
@@ 6 Experiments
The goal of this section is to verify empirically that existing state-of-the-art methods for addressing sub-population shifts benefit from using optimized group weights. We use benchmark classification datasets for studying sub-population shifts: two from computer vision (Waterbirds, CelebA) and one from natural language processing (MultiNLI). Groups are defined via a combination of the target label and a spurious attribute. The training and validation set are randomly split. For Waterbirds, this is different from previous implementations, where the validation set is balanced in terms of the groups (Sagawa et al., 2020a). We deviate from this, since it is more realistic that the training and validation sets are drawn from the same distribution. The distribution shift is established by measuring the generalization performance on a test set for which all groups have an equal weight. For details on the datasets, including group sizes, see Appendix E.1.
@@
Furthermore, we find that optimizing weights often also improves the metric for which the weights were not optimized (worst-group accuracy for GW-ERM, SUBG and DFR; weighted average accuracy for GDRO and JTT). The only exception is SUBG on CelebA. In 10 cases, there is an improvement greater than 1%, with a maximum of 4.03%. This result was not necessarily expected, since weighted average accuracy and group robustness are often opposed objectives (Agarwal et al., 2018; Martinez et al., 2021a). See appendix C for numerical details of the results.
@@
We test this for GW-ERM on CelebA in Figure 3, where we show the effect of optimized weights using only fractions of the original training and validation set (while keeping the ratio of the sizes the same). The improvement due to optimizing weights peaks when around 10% of the original data is used. Note that for smaller fractions of the data (5%) the effect decreases again. This is likely due to the validation set becoming (very) small, increasing the chance of overfitting on this data.
@@ 7 Conclusion and Discussion
This paper has shown that the standard choice of weights for importance weighting in the presence of sub-population shifts is often not optimal, due to the finite sample size of the training data. We have provided an estimation procedure for the optimal weights, which generally improves existing importance weighting methods in the case of last-layer retraining of DNNs.
@@ Appendix A Calculation of the hyper-gradient
$\frac{\partial\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}{\partial\bm{p}}=-\left[\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{\theta}^{T}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}\right]^{-1}\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{p}^{T}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}.$ (15)
@@
where we used that $p_{G}=1-\sum_{g^{\prime}=1}^{G}p_{g^{\prime}}$ due to normalization. Then, in order to determine $\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{p}^{T}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}$, we can calculate the derivative with respect to the estimated parameters for the novel loss function in Equation 17 for each $p_{g}$.
Subsampling large groups: here, we are interested in the hyper-gradient for the hyperparameter $\bm{v}$ instead of $\bm{p}$, as defined in Section 5.2. The parameters $\hat{\bm{\theta}}_{n^{\prime}}(\bm{v})$ are not determined via a weighted loss, but rather by minimizing the training loss on a subsample of the data. However, the expected training loss can be rewritten as a weighted loss. To mimic the procedure of subsampling without replacement, we say that for each $g=1,\ldots,G$ the random variables $s^{{g}}_{{i}}$ are identically distributed $\mathrm{Bern}(v_{g})$ random variable with constraint
@@
Note the difference with Equation 17, because the $\bm{v}$ are not normalized. To determine $\left.\frac{\partial^{2}\tilde{h}(\bm{\theta},\bm{v})}{\partial\bm{\theta}\partial\bm{v}^{T}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{v})}$, we can (again) calculate the derivative with respect to the estimated parameters for the novel loss function in Equation 21 for each $v_{g}$. In addition to the change in the calculation of the hyper-gradient, the only change to Algorithm 1 is that we do not normalize the weights $\bm{v}$.
@@ Appendix B Theoretical analysis for linear regression
  • In Section B.1 we give a description how $\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[(y-\bm{x}^{T}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right]$ can be decomposed into a bias and variance term in the case of linear regression, and data stemming from $G$ groups.

### v2

@@ 1 Introduction
Secondly, we introduce an estimation procedure for the optimal weights via a bi-level optimization problem, where the weights and the model parameters are optimized simultaneously. In the context of last-layer retraining of DNNs in the presence of sub-population shifts, we show empirically, for benchmark vision and natural language processing datasets, that existing state-of-the-art importance weighting methods (see Section 2) can be improved significantly by optimizing for the weights. We find that both the weighted average and worst-group accuracy generally improve, and that optimized weights increase robustness against the choice of hyperparameters for training. We show that the effect of optimizing weights is larger when the limited size of the training sample becomes pressing, namely in the case of a small total sample size or when the minority groups are small. We provide an implementation of this procedure as an open-source package (link).
@@ 2 Related work
Sub-population shifts: it has been widely documented that ML models frequently face deteriorating generalization performance due to sub-population shifts (Cai et al., 2021; Koh et al., 2021; Yang et al., 2023). There exist different possible definitions of the groups for which the distribution shift occurs. In the case of a spurious correlation, groups are commonly defined by a combination of a label and a non-causally related attribute (Joshi et al., 2022; Makar et al., 2022). Alternatively, groups can be defined exclusively via an attribute (Martinez et al., 2021b) or, in the case of class imbalance, exclusively via a label (Liu et al., 2019). While our experiments consider spurious correlations, the methodology also applies to alternative group definitions. Generalizing to groups that do not occur in the training data (Santurkar et al., 2020) is beyond the scope of this paper.
@@
Estimation of optimal weights: a commonly noted downside of importance weighting is that it reduces the effective sample size available for model training (Martino et al., 2017; Schwartz & Stanovsky, 2022). In light of this, Cui et al. (2019) and Wen et al. (2020) suggest that the weight given to a particular label should be based on the number of available datapoints. Similar to our approach, Ren et al. (2018) aim to learn weights for neural networks given a validation set.
@@ 3 Problem Setting
Consider random variables $(y,\bm{x}),$ where $y\in\mathcal{Y}$ is the target variable and $\bm{x}\in\mathcal{X}$ represents the input features. Suppose we are dealing with a family of models $\left\{f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Y}\,|\,\bm{\theta}\in\Theta\right\}$ and we aim to find a model that minimizes $\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[\mathcal{L}(y,f_{\bm{\theta}}(\bm{x}))],$ where $\mathcal{L}(\cdot,\cdot)$ is some pre-defined loss function and $\mathcal{P}_{\mathrm{te}}$ is a test distribution of interest. Moreover, we have access to a dataset $D_{\mathrm{tr}}^{n}=\{y_{{i}},\bm{x}_{i}\}^{n}_{i=1},$ which is an i.i.d. sample drawn from a training distribution $\mathcal{P}_{\mathrm{tr}}$ that is different from the test distribution. This distribution shift can be problematic, as the minimum of an empirical loss for the training data does typically not coincide with the minimum expected loss with respect to the test distribution.
@@
If we choose the weights $\bm{w}$ to be equal to the likelihood ratios $\bm{r},$ then by the law of large numbers $\mathcal{L}_{\mathrm{train}}^{n}(\bm{\theta},\bm{r})$ converges in the limit of $n\rightarrow\infty$ to the expectations in Equation 2. As a consequence, $\hat{\bm{\theta}}_{n}(\bm{r})$ minimizes for sample size tending to infinity the expected loss with respect to the test distribution, $\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[\mathcal{L}(y,f_{\bm{\theta}}(\bm{x})].$ However, in practice one works with a finite amount of training data and the expectation might be poorly approximated by its sample analogue. Therefore, the optimal choice $\bm{w}^{*}_{n}$ for the weights should be defined as the one that leads to a “best” parameter estimate $\hat{\bm{\theta}}_{n}(\bm{w}),$ in the sense that it minimizes the expected loss with respect to the test distribution,
@@ 4 Theoretical analysis of linear regression
We consider a DGP $y=\bm{x}^{\mathrm{T}}\bm{\beta}^{g}+\varepsilon,$ where $y\in\mathbb{R}$ is the target variable, $\bm{x}=(1,\tilde{\bm{x}}^{\mathrm{T}})^{\mathrm{T}}$ represents the intercept and the feature vector $\tilde{\bm{x}}\in\mathbb{R}^{d},$ and $\varepsilon\in\mathbb{R}$ is the noise. We take $\tilde{\bm{x}}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ and the noise to have zero mean, variance $\sigma^{2}$ and to be uncorrelated with $\bm{x}.$ The population consists of two groups labeled by $g\in\{0,1\}.$ Group membership determines the value of the intercept, $\bm{\beta}^{g}=(a_{g},\bm{\beta}^{\mathrm{T}})^{\mathrm{T}}$ for $g=0,1.$ Note that the slope coefficients are common among the two groups. The sub-population shift between training and test data is defined by a change in the marginal distribution of the group label, $p_{\mathrm{tr}}:=\mathbb{P}_{\mathrm{tr}}(g=1)\neq\mathbb{P}_{\mathrm{te}}(g=1)=:p_{\mathrm{te}}.$
@@
$\hat{\bm{\beta}}_{n}(p)=(\bm{X}^{\mathrm{T}}\bm{W}\bm{X})^{-1}\bm{X}^{\mathrm{T}}\bm{W}\bm{y},\quad\text{with }\bm{W}=\mathrm{diag}(w_{1},...,w_{n}),\hskip 2.84544ptw_{i}=\begin{cases}\frac{p}{p_{\mathrm{tr}}}&g_{i}=1,\\ \frac{1-p}{1-p_{\mathrm{tr}}}&g_{i}=0.\end{cases}$ (7)
@@
$\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}\left[(y-\bm{x}^{\mathrm{T}}\hat{\bm{\beta}}_{n}(\bm{p}))^{2}\middle|D_{\mathrm{tr}}^{n}\right]\right]\approx\mathcal{B}^{2}(p_{\mathrm{te}},p,a_{1},a_{0})+\mathcal{V}(\sigma^{2},p,p_{\mathrm{tr}},n,d)+\sigma^{2},$ (8a)
@@
This is the optimal trade-off between minimizing the bias and variance. In the limit $n\rightarrow\infty$ the variance disappears and the optimal $p$ is indeed equal to the heuristic choice $p_{\mathrm{te}}.$ However, for finite $n$ it can differ considerably. This is illustrated in Figure 1. Its impact on the expected loss is shown in Figure 5 in Appendix B.
@@ 5.1 Group-Weighted Empirical Risk Minimization
We solve the bi-level optimization problem iteratively. We find $\hat{\bm{p}}^{*}_{n}$ in Equation 10 by means of exponentiated gradient descent (Nemirovsky & Yudin, 1983; Kivinen & Warmuth, 1997), where the parameter estimates $\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})$ are kept fixed. After each weight update, we re-calibrate the parameter estimates of Equation 1. Although convergence to the optimal group weights is not guaranteed due to potential non-convexity, gradient-based optimization of hyperparameters has previously been applied successfully (Pedregosa, 2016). Details of the optimization procedure are provided in Algorithm 1.
@@
Split the data into a training set of size $n^{\prime}$ and a validation set. Set $u_{0,g}=0$ for all $g\in\mathcal{G}$.
@@ 5.3 Group Distributional Robust Optimization
where, as before, $\mathcal{P}_{G}$ is the $G$-dimensional probability simplex. The inner expectation of Equation 13 is estimated with a validation set that is withheld from the data used to train the model parameters. The bi-level optimization procedure now consists of updating three quantities per update step: the model parameters $\bm{\theta},$ the group weights $\bm{p}$ and the loss weights $\bm{q}.$ See Algorithm 2 in Appendix D for details.
@@ 6 Experiments
The goal of this section is to verify empirically that existing state-of-the-art methods for addressing sub-population shifts benefit from using optimized group weights. We use benchmark classification datasets for studying sub-population shifts: two from computer vision (Waterbirds, CelebA) and one from natural language processing (MultiNLI). Groups are defined via a combination of the target label and a spurious attribute. The training and validation set are randomly split. For Waterbirds, this is different from previous implementations, where the validation set is balanced in terms of the groups (Sagawa et al., 2020a). We deviate from this, since it is more realistic that the training and validation sets are drawn from the same distribution. The distribution shift is established by measuring the generalization performance on a test set for which the groups have equal weights. For details on the datasets, including group sizes, see Appendix E.1.
@@
Furthermore, we find that optimizing weights often also improves the metric for which the weights were not optimized (worst-group accuracy for GW-ERM, SUBG and DFR; weighted average accuracy for GDRO and JTT). The only exception is SUBG on CelebA. In 10 cases, there is an improvement by more than 1%, with a maximum of 4.03%. This result was not necessarily expected, since weighted average accuracy and group robustness are often opposed objectives (Agarwal et al., 2018; Martinez et al., 2021a). See appendix C for numerical details of the results.
@@
We test this for GW-ERM on CelebA in Figure 3, where we show the effect of optimized weights using only fractions of the original training and validation set (while keeping the ratio of the sizes the same). The improvement due to optimizing weights peaks when around 10% of the original data is used. Note that for smaller fractions of the data (5%) the effect decreases again. This is likely due to the validation set becoming (very) small, increasing the risk of overfitting on this data.
@@ 7 Conclusion and Discussion
In this paper we have shown that the standard choice of weights for importance weighting in the presence of sub-population shifts is often not optimal, due to the finite sample size of the training data. We have provided an estimation procedure for the optimal weights, which generally improves existing importance weighting methods in the case of last-layer retraining of DNNs.
@@ Appendix A Calculation of the hyper-gradient
$\frac{\partial\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}{\partial\bm{p}}=-\left[\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{\theta}^{\mathrm{T}}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}\right]^{-1}\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{p}^{\mathrm{T}}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}.$ (15)
@@
where we used that $p_{G}=1-\sum_{g^{\prime}=1}^{G}p_{g^{\prime}}$ due to normalization. Then, in order to determine $\left.\frac{\partial^{2}h(\bm{\theta},\bm{p})}{\partial\bm{\theta}\partial\bm{p}^{\mathrm{T}}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{p})}$, we can calculate the derivative with respect to the estimated parameters for the novel loss function in Equation 17 for each $p_{g}$.
Subsampling large groups: here, we are interested in the hyper-gradient for the hyperparameter $\bm{v}$ instead of $\bm{p}$, as defined in Section 5.2. The parameters $\hat{\bm{\theta}}_{n^{\prime}}(\bm{v})$ are not determined via a weighted loss, but rather by minimizing the training loss on a subsample of the data. However, the expected training loss can be rewritten as a weighted loss. To mimic the procedure of subsampling without replacement, we say that for each $g=1,\ldots,G$ the random variables $s^{{g}}_{{i}}$ are identically distributed $\mathrm{Bern}(v_{g})$ random variables with constraint
@@
Note the difference with Equation 17, because the $\bm{v}$ are not normalized. To determine $\left.\frac{\partial^{2}\tilde{h}(\bm{\theta},\bm{v})}{\partial\bm{\theta}\partial\bm{v}^{\mathrm{T}}}\right|_{\bm{\theta}=\hat{\bm{\theta}}_{n^{\prime}}(\bm{v})}$, we can (again) calculate the derivative with respect to the estimated parameters for the novel loss function in Equation 21 for each $v_{g}$. In addition to the change in the calculation of the hyper-gradient, the only change to Algorithm 1 is that we do not normalize the weights $\bm{v}$.
@@ Appendix B Theoretical analysis for linear regression
  • In Section B.1 we give a description of how $\mathbb{E}_{D_{\mathrm{tr}}^{n}\sim\mathcal{P}_{\mathrm{tr}}}\!\left[\mathbb{E}_{(y,\bm{x})\sim\mathcal{P}_{\mathrm{te}}}[(y-\bm{x}^{\mathrm{T}}\hat{\bm{\beta}}(\bm{w}))^{2}|D_{\mathrm{tr}}]\right]$ can be decomposed into a bias and variance term in the case of linear regression, and data stemming from $G$ groups.